OpenStudy (rockthered1776):

Solve for v (5v-1)(5-v)=0 Just need someone to walk me through it

OpenStudy (retireed):

There are two values of "v" that will cause the equation to equal zero. So you can solve for them separately by setting each "piece" of the left side of the equation equal to zero and solving them separately. (5v-1)(5-v)=0 can be separated into (5v-1)=0 and (5-v)=0 Solve each of the simpler pieces for v. Do you understand so far?

OpenStudy (rockthered1776):

yes, i believe I follow so far... for problem 1 Id move the 1 to the other side and divide by 5 for an answer of 1/5... problem 2 id move the 5 over multiply both sides by -1 and get v=5 correct? @retirEEd

OpenStudy (rockthered1776):

Think I made this problem much more complicated than it had to be

OpenStudy (retireed):

That look correct to me. You have a good grasp of the algebra.

OpenStudy (rockthered1776):

Appreciate the help.. just needed a refresher on how to go about the problem. Thank you for your time. @retirEEd

OpenStudy (retireed):

yw